### v4

@@ 1.1 Random projections in statistics
Random-projection (RP) techniques have become a cornerstone in high-dimensional statistical analysis. Over the last two decades, the theoretical and applied developments of RP have flourished. Early applications emphasized computational efficiency and scalability, particularly for clustering [8], matrix approximations, [70, 41], and kernel learning [65]. The work presented in [1] introduced more efficient sparse and structured random matrices, which proved crucial for large-scale implementations.
@@ 1.1.1 RP via the Johnson–Lindenstrauss lemma
Lemma (Johnson–Lindenstrauss lemma, [46]). Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
@@
Recent studies continue to build on the Johnson–Lindenstrauss lemma, investigating its application in modern contexts, for example, as a method for dimensionality reduction [22], particularly in machine-learning applications [30], in high-dimensional clustering [2], and nearest neighbor search [49]. These studies demonstrate how RP can preserve important properties in high-dimensional spaces while dramatically reducing the computational complexity of various algorithms.
@@ 1.1.2 RP via the Cramér–Wold theorem
Theorem 2 (Cuesta–Fraiman–Ransford [18, Corollary 3.2]).
@@
In contrast to mainstream applications of random projection, such as kernel approximation [65] or compressed sensing [11], the approach developed in [17, 18] is distinguished by its nonparametric nature and inferential focus. Rather than focusing on data transformations that preserve structural relationships for machine learning, the latter work is more oriented to preserving the geometric properties of the data for inferential purposes. This distinction in focus underscores the role of RP in nonparametric inference, as it provides an intuitive tool for deriving statistical conclusions from data without imposing stringent parametric assumptions.
@@
Theorem 3 (Fraiman–Moreno–Ransford [35]).
@@ 1.1.3 On the pros and cons of RP-methods
As modern datasets continue to grow in complexity and dimensionality, the statistical insight provided by projection-based depth and testing becomes increasingly relevant. It offers a complement—and at times a counterpoint—to the more engineering-oriented uses of RP found in neural networks [14], sketching algorithms [50], or NLP embeddings [77]. The blend of theoretical rigor and practical robustness in the works of Cuesta, Fraiman, Ransford and their collaborators provides tools that are especially suitable for applied statisticians working under real-world constraints, such as noisy data or nonstandard distributions.
@@ 2.1 Mixtures
Mixture models are fundamental tools in statistics and machine learning, providing a flexible framework for representing heterogeneous data as combinations of simpler component distributions. Since the pioneering work of Pearson [61], they have been extensively used in tasks such as clustering [74], density estimation [48], and the modeling of latent structures [59]. Finite mixture models have also found applications across a broad range of disciplines, including astronomy, biology, genetics, economics, social sciences, and engineering [52].
@@
Gaussian-mixture models have proven to be highly effective in modeling various real-world data. For a comprehensive examination of their properties, we refer to Titterington et al. [75]. These models are flexible and general, with relevant applications documented in numerous fields, including density estimation, machine learning, and clustering. As noted in [13], significant applications include Pearson’s work on modeling crab data [61], Roeder’s studies [69], and the research by Schork, Allison and Thiel in genetics [71], as well as Everitt’s investigations involving schizophrenia patients [29], among many others.
@@ 2.2 A Cramér–Wold Theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic Theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [34]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [35, Theorems 1 and 2], which we will now summarize.
@@
We now recall the Cramér–Wold Theorem for Gaussian mixtures [34, Theorem 4.1].
Theorem 4.
@@
We also recall an analogue of Theorem 4 for mixtures of multivariate $t$-distributions, given in [34, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 5.
@@ 2.3 Strong sm-uniqueness sets
Theorems 4 and 5 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [34, Theorem 4.3], which also provides a realistic method for generating them.
@@
Definition 1 (sm-uniqueness and strong sm-uniqueness sets [35, 34]).
@@ 3.1 Introduction
There is a large literature on this subject. See for instance the well-known work in [54], the recent work [27] and the references therein. The article [54] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}-f_{j}|\geq\epsilon$. In [27], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before [34, Theorem 4.1] below.
@@ 3.2 Description of the algorithm
5:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve, for $j=1,\dots,n$,
@@
In Algorithm 1, the parameters are initialized separately in each random one-dimensional projection and the main driver of the initialization is a $k$-means split. Specifically, for a randomly sampled direction $u$ (normalized to unit length), the data are projected as $Y_{i}=u^{\top}X_{i}$ and we run $k$-means with two clusters on $\{Y_{i}\}_{i=1}^{N}$. The resulting two clusters provide a data-driven starting point: the initial projected locations are set to the within-cluster sample means, the projected scales to the within-cluster sample variances (truncated below by a small constant for numerical stability), and the initial mixing weight is taken as the empirical cluster proportion. To avoid poor starts, we only keep projections where both clusters contain at least a minimum number of observations and the two $k$-means centroids are sufficiently separated relative to $\mathrm{sd}(Y)$. After ordering components by increasing projected mean, the $k$-means-based initialization is optionally refined by a single soft-EM update for the mixing weight, and then passed to the GEL estimation step in that projection.
@@
for all $j=1,2,\dots,n$.
Alternatively, in the spirit of the proposed median of means, see for instance [25] and the references therein, one can proceed as as follows. We split the sample into $L$ disjoint subsamples. For each subsample and $j$ we calculate $\widehat{\mu}_{j,\ell}$ and $\widehat{\Sigma}_{j,\ell}$, $j=1,\dots n$, $\ell=1,\ldots,L$ according to equations (4) and (5), and we take the median of the corresponding $L$ values obtained for each $j$. Then we proceed as before.
(iii) The $\operatorname*{argmin}$ in (4) can be easily solved using least squares. To solve the constrained optimization problem of estimating covariance matrices under positive semi-definiteness, see (5), we apply a primal-dual interior-point method (IPM). This approach is well-suited for semidefinite programming (SDP), as it ensures convergence to the global optimum within the cone of symmetric positive semidefinite matrices. The method combines Newtonian steps with a logarithmic barrier function to keep iterates in the feasible region, which in our case corresponds to the set $\mathbb{S}^{+,d}$. This strategy is particularly appropriate given the convexity of our loss function and the structure of the constraint. For background and theoretical guarantees, see [81, 57, 68].
Since $f(\Sigma)=\Bigl(\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr)^{2}$ is convex and continuous on $\mathbb{S}^{+,d}$, the optimization problem admits a unique solution. Consequently, the penalized objective function used in the IPM Algorithm, $f_{t}(\Sigma)=tf(\Sigma)-\log\det(\Sigma)$, attains a unique minimum over $\mathbb{S}^{+,d}$; see, e.g., [7, 10].
@@ 3.3 A consistency result
the estimators $\widehat{\Lambda}_{(1)},\widehat{\mu}_{j},\widehat{\Sigma}_{j}$ for $j=1,\dots,n$ are strongly consistent.
@@
First observe that, from the results in [76] and [82] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [76, Proposition 1], where strong consistency and asymptotic normality are established.
For the next step, using the value of $\widehat{\Lambda}_{(1)}$, we consider the estimators of $\widehat{\mu}_{j}$ and $\widehat{\Sigma}_{j}$ for $j=1,\dots n$ given by the least-squares equations (4) and (5). Since
@@
for the true values of $\mu_{j}$ and $\Sigma_{j}~(j=1,\dots n)$, we have
@@
for all $j=1,\dots,n$. This concludes the proof. ∎
The only assumption in our Theorem 7 is that $I_{n}(t)$ can be differentiated under the integral sign. As mentioned in [43, p. 258], the only random variables excluded are those with characteristic functions vanishing outside an interval depending on the parameter $\theta$ and those for which $\psi$ is not differentiable with respect to $\theta$. These conditions are fulfilled for the univariate mixtures that we have under consideration. Indeed, the differentiation condition follows for instance from [63].
@@
If one is only interested on D(F,F^)D($\widehat{F}$), then in [55, Corollary 2] there is stated a similar result without the restriction that $F$ be $\epsilon$-statistically learnable.
@@ 4.2 Description of the algorithm
Suppose that we want to compare the results obtained from two clustering procedures, both applied to the same set of variables. We consider two independent samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$, with distributions $P$ and $Q$ respectively, both being mixtures of $m$ multivariate Gaussian or $t$-distributions.
@@
As $\min(\ell,r)\to\infty$, the asymptotic distribution of $D_{k}$ satisfies
$\lim_{\min(\ell,r)\to\infty}\Pr(D_{k}\leq t)=\lim_{\min(\ell,r)\to\infty}\Big[\Pr(\mathrm{KS}(1)\leq t)\Big]^{k},$
which provides the basis for an asymptotic test. A drawback is the loss of the distribution-free property, as the distribution of $D_{k}$ depends on the covariance structure of the underlying processes.
The overall scheme is summarized in Algorithm 2.
@@
1:Integers $m,d$; samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$, $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)+1$
@@
4:  Project samples: $\mathcal{M}_{1i}=\{\langle u_{i},X_{j}\rangle\}_{j=1}^{\ell}$,  $\mathcal{M}_{2i}=\{\langle u_{i},Y_{j}\rangle\}_{j=1}^{r}$
@@
8:return $\mathrm{D}_{k}\leftarrow\max_{i=1,\dots,k}\mathrm{KS}(i)$
@@
(i) An alternative measure is the average
$\mathrm{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i),$
where $\mathrm{KS}(i)$ is the KS distance between the empirical distributions of $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$. If $F_{i}$ and $G_{i}$ are the true projected distributions, then by the triangle inequality,
$0\leq\mathrm{KS}(i)\leq\mathrm{KS}(F_{\mathcal{M}_{1i}},F_{i})+\mathrm{KS}(F_{i},G_{i})+\mathrm{KS}(G_{\mathcal{M}_{2i}},G_{i}),$
which implies
$\lim_{\min(\ell,r)\to\infty}\mathrm{MA}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(F_{i},G_{i}).$
Theorems 4 and 5 show that if this limit is zero, then $P=Q$. However, the asymptotic distribution of $\mathrm{MA}_{k}$ is unknown, since projections onto different directions are only conditionally independent.
(ii) If the sample sizes $\ell$ and $r$ are large, one may split them into $k$ disjoint subsamples and use different subsamples for each direction $u_{i}$. In this case, $\mathrm{MA}_{k}$ (or $D_{k}$) becomes an average (or maximum) of $k$ independent one-dimensional KS statistics, which is distribution-free. This strategy, however, comes at the cost of reduced statistical power.
(iii) Pre-whitening via $\Sigma^{-1/2}$ is a convenient option to mitigate scale (and unit-of-measure) effects when the practitioner wishes to control for them; in that case, one may apply a whitening transform to the data before computing the projected KS aggregations.

### v5

@@ 1.1 Random projections in statistics
Random-projection (RP) techniques have become a cornerstone in high-dimensional statistical analysis. Over the last two decades, the theoretical and applied developments of RP have flourished. Early applications emphasized computational efficiency and scalability, particularly for clustering and matrix approximations [6, 35]. The work presented in [1] introduced more efficient sparse and structured random matrices, which proved crucial for large-scale implementations.
@@ 1.1.1 RP via the Johnson–Lindenstrauss lemma
Lemma ([40]).
Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
@@
Recent studies continue to build on the Johnson–Lindenstrauss lemma, investigating its application in modern contexts, for example, as a method for dimensionality reduction [18], particularly in machine-learning applications, in high-dimensional clustering, and nearest neighbor search. These studies demonstrate how RP can preserve important properties in high-dimensional spaces while dramatically reducing the computational complexity of various algorithms.
@@ 1.1.2 RP via the Cramér–Wold theorem
Theorem 2 ([14, Corollary 3.2]).
@@
In contrast to mainstream applications of random projection, such as kernel approximation or compressed sensing, the approach developed in [13, 14] is distinguished by its nonparametric nature and inferential focus. Rather than focusing on data transformations that preserve structural relationships for machine learning, the latter work is more oriented to preserving the geometric properties of the data for inferential purposes. This distinction in focus underscores the role of RP in nonparametric inference, as it provides an intuitive tool for deriving statistical conclusions from data without imposing stringent parametric assumptions.
@@
Theorem 3 ([29, Theorem 1 and Proposition 1]).
@@ 1.1.3 On the pros and cons of RP-methods
As modern datasets continue to grow in complexity and dimensionality, the statistical insight provided by projection-based depth and testing becomes increasingly relevant. It offers a complement—and at times a counterpoint—to the more engineering-oriented uses of RP found in neural networks, sketching algorithms, or NLP embeddings. The blend of theoretical rigor and practical robustness in the works of Cuesta, Fraiman, Ransford and their collaborators provides tools that are especially suitable for applied statisticians working under real-world constraints, such as noisy data or nonstandard distributions.
@@ 2.1 Mixtures
Mixture models are fundamental tools in statistics and machine learning, providing a flexible framework for representing heterogeneous data as combinations of simpler component distributions. Since the pioneering work of Pearson [49], finite mixture models have been widely used for clustering, density estimation, and latent structure modeling [43, 42]. They have also found applications across many disciplines.
@@
Gaussian-mixture models have proven to be highly effective in modeling various real-world data. For a comprehensive examination of their properties, we refer to Titterington et al. [58]. These models are flexible and general, with relevant applications documented in numerous fields, including density estimation, machine learning, and clustering. As noted in [10], significant applications include Pearson’s early work on modeling crab data [49], among many others.
@@ 2.2 A Cramér–Wold Theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic Theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [28]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [29, Theorems 1 and 2], which we will now summarize.
@@
We now recall the Cramér–Wold Theorem for Gaussian mixtures from [28].
Theorem 4 ([28, Theorem 4.1]).
@@
We also recall an analogue of Theorem 4 for mixtures of multivariate $t$-distributions, given in [28, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 5 ([28, Theorem 4.2]).
@@ 2.3 Strong sm-uniqueness sets
Theorems 4 and 5 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [28, Theorem 4.3], which also provides a realistic method for generating them.
@@
Definition 1 ([29, 28]).
@@ 3.1 Introduction
There is a large literature on this subject. See for instance the well-known work in [45], the recent work [23] and the references therein. The article [45] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}-f_{j}|\geq\epsilon$. In [23], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before [28, Theorem 4.1] below.
@@ 3.2 Description of the algorithm
5:Step 2: Fix $\widehat{\Lambda}^{(1)}$ and re-estimate $\{(\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)})\}_{r=1}^{k}$ (Appendix A).
6:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve, for $j=1,\dots,m$,
@@
In Algorithm 1, the parameters are initialized separately in each random one-dimensional projection and the main driver of the initialization is a $k$-means split, (with $k=m$). Specifically, for a randomly sampled direction $u$ (normalized to unit length), the data are projected as $Y_{i}=u^{\top}X_{i}$ and we run $k$-means with $m$ clusters on $\{Y_{i}\}_{i=1}^{N}$. The resulting $m$ clusters provide a data-driven starting point: the initial projected locations are set to the within-cluster sample means, the projected scales to the within-cluster sample variances (truncated below by a small constant for numerical stability), and the initial mixing weight is taken as the empirical cluster proportion. To avoid poor starts, we only keep projections where both clusters contain at least a minimum number of observations and the two $k$-means centroids are sufficiently separated relative to $\mathrm{sd}(Y)$. After ordering components by increasing projected mean, the $k$-means-based initialization is optionally refined by a single soft-EM update for the mixing weight, and then passed to the GEL estimation step in that projection.
@@
for all $j=1,2,\dots,m$.
Alternatively, in the spirit of the proposed median of means, see for instance [21] and the references therein, one can proceed as as follows. We split the sample into $L$ disjoint subsamples. For each subsample and $j$ we calculate $\widehat{\mu}_{j,\ell}$ and $\widehat{\Sigma}_{j,\ell}$, $j=1,\dots m$, $\ell=1,\ldots,L$ according to equations (4) and (5), and we take the median of the corresponding $L$ values obtained for each $j$. Then we proceed as before.
(iii) The $\operatorname*{argmin}$ in (4) can be easily solved using least squares. To solve the constrained optimization problem of estimating covariance matrices under positive semi-definiteness, see (5), we apply a primal-dual interior-point method (IPM). This approach is well-suited for semidefinite programming (SDP), as it ensures convergence to the global optimum within the cone of symmetric positive semidefinite matrices. The method combines Newtonian steps with a logarithmic barrier function to keep iterates in the feasible region, which in our case corresponds to the set $\mathbb{S}^{+,d}$. This strategy is particularly appropriate given the convexity of our loss function and the structure of the constraint. For background and theoretical guarantees, see [55].
Since $f(\Sigma)=\Bigl(\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr)^{2}$ is convex and continuous on $\mathbb{S}^{+,d}$, the optimization problem admits a unique solution. Consequently, the penalized objective function used in the IPM Algorithm, $f_{t}(\Sigma)=tf(\Sigma)-\log\det(\Sigma)$, attains a unique minimum over $\mathbb{S}^{+,d}$; see, e.g., [8].
@@ 3.3 A consistency result
the estimators $\widehat{\Lambda}_{(1)},\widehat{\mu}_{j},\widehat{\Sigma}_{j}$ for $j=1,\dots,m$ are strongly consistent.
@@
First observe that, from the results in [59] and [63] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [59, Proposition 1], where strong consistency and asymptotic normality are established.
For the next step, using the value of $\widehat{\Lambda}_{(1)}$, we consider the estimators of $\widehat{\mu}_{j}$ and $\widehat{\Sigma}_{j}$ for $j=1,\dots m$ given by the least-squares equations (4) and (5). Since
@@
for the true values of $\mu_{j}$ and $\Sigma_{j}~(j=1,\dots m)$, we have
@@
for all $j=1,\dots,m$. This concludes the proof. ∎
The only assumption in our Theorem 7 is that $I_{n}(t)$ can be differentiated under the integral sign. As mentioned in [37, p. 258], the only random variables excluded are those with characteristic functions vanishing outside an interval depending on the parameter $\theta$ and those for which $\psi$ is not differentiable with respect to $\theta$. These conditions are fulfilled for the univariate mixtures that we have under consideration. Indeed, the differentiation condition follows for instance from [51].
@@
If one is only interested on D(F,F^)D($\widehat{F}$), then in [45, Corollary 2] there is stated a similar result without the restriction that $F$ be $\epsilon$-statistically learnable.
@@ 4.2 Description of the algorithm
Suppose we wish to compare the results obtained from two clustering procedures. We consider two independent samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$, with distributions $P$ and $Q$ respectively, both being mixtures of $m$ multivariate Gaussian or $t$-distributions.
@@
An alternative measure is the average
$\mathrm{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i),$
where $\mathrm{KS}(i)$ is the KS distance between the empirical distributions of $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$.
Theorems 4 and 5 show that if this limit is zero, then $P=Q$. The asymptotic distributions of $\mathrm{MA}_{k}$ and $D_{k}$ are unknown, since projections onto different directions are only conditionally independent. However, we can use the bootstrap method to proceed in practice.
Alternatively if the sample sizes $\ell$ and $r$ are large, we split them into $k$ disjoint subsamples $\aleph_{split}:=(\aleph_{11},\aleph_{21}),\ldots((\aleph_{1k},\aleph_{2k}))$, and use different subsamples for each direction $u_{i}$. In this case, $\mathrm{MA}_{k}$ (or $D_{k}$) becomes an average (or maximum) of $k$ independent one-dimensional KS statistics, which is distribution-free. This strategy, however, comes at the cost of reduced statistical power.
The overall scheme is summarized in Algorithm 2 for $k$ disjoint subsamples; an analogous scheme can be implemented for the bootstrap method.
@@
1:Integers $m,d$; subsamples $\aleph_{split}$ ; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)+1$
@@
4:  Project subsamples: $\mathcal{M}_{1i}=\{\langle u_{i},\aleph_{1i}\rangle\}$,  $\mathcal{M}_{2i}=\{\langle u_{i},\aleph_{2i}\rangle\}$
@@
8:return $\mathrm{D}_{k}\leftarrow\max_{i=1,\dots,k}\mathrm{KS}(i)$ or $\mathrm{M}_{k}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i)$
@@
  1. 1.
    Pre-whitening via $\Sigma^{-1/2}$ is a convenient option to mitigate scale (and unit-of-measure) effects when the practitioner wishes to control for them; in that case, one may apply a whitening transform to the data before computing the projected KS aggregations.
  2. 2.
    In the one-sample setting with two clustering procedures, we start from a single sample $\aleph=\{X_{1},\dots,X_{N}\}\subset\mathbb{R}^{d}$. We split $\aleph$ into two disjoint subsamples $\aleph_{1}$ and $\aleph_{2}$, and apply a different clustering procedure to each subsample (e.g., GMMs fitted under different prior specifications). Note that if the data are mapped through a linear transformation (in particular, a one-dimensional projection) while the GMM is fitted in the original space, then the induced distribution on the projection subspace is again a Gaussian mixture, with mean vectors and covariance matrices obtained by applying the same linear map to the fitted parameters. Consequently, for each random direction we may fit a univariate Gaussian mixture to the projected data, using mixture weights given by the cluster proportions, and then compute $D_{k}$ or $\mathrm{MA}_{k}$.